OpenStudy (anonymous):

Solve using substitution 3x - 9y = 3 6x - 3y = -24 Solve using elimination y - 1/2x = 6 2x + 6y = 19

OpenStudy (mathstudent55):

For the first system, start by dividing the first equation by 3 and the second equation also by 3.

OpenStudy (mathstudent55):

What do you get for the first equation after dividing by 3?

OpenStudy (the_fizicx99):

Convert one into y = mx + b then plug in that for y into the other eqaution 3x - 9y = 3 6x - 3y = -24 6x - 3y = -24 -6x -6x -3y = -6x -24 /-3 /-3 /-3 y = 2x + 8 3x - 9(2x + 8) = 3 Can you solve from there?